OpenStudy (anonymous):

How to write 3x3 matrix?

OpenStudy (anonymous):

\( \begin{array}{ccc} a_{1,1} & a_{1,2} & a_{1,3} \\ a_{2,1} & a_{2,2} & a_{2,3} \\ a_{3,1} & a_{3,2} & a_{3,3} \\ \end{array} \)

OpenStudy (unklerhaukus):

\[\left(\begin{array}{ccc} a_{1,1} & a_{1,2} & a_{1,3} \\ a_{2,1} & a_{2,2} & a_{2,3} \\ a_{3,1} & a_{3,2} & a_{3,3} \\ \end{array}\right)\]

OpenStudy (turingtest):

\left[\.begin{matrix}\end{matrix}\right] (take out the . before the word "begin") sets you up for any sized matrix: \left[\.begin{matrix}1&2&3\\4&5&6\\7&8&9\end{matrix}\right] \[\left[\begin{matrix}1&2&3\\4&5&6\\7&8&9\end{matrix}\right]\] notice the & lets you put an entry in the next column, and \\ lest you make entries in the next row you can make any size or shape matrix with this also if you put \left(\right) you get the borders that @UnkleRhaukus wrote you can also put \left|\right| to make it like a determinant\[\left|\begin{matrix}1&2&3\\4&5&6\\7&8&9\end{matrix}\right|\]or \left.\right. so that nothing appears to border the matrix\[\left.\begin{matrix}1&2&3\\4&5&6\\7&8&9\end{matrix}\right.\]

OpenStudy (unklerhaukus):

\.begin{array}\ 1&2&3\\4&5&6\\7&8&9\end{array}\[\begin{array}\ 1&2&3\\4&5&6\\7&8&9\end{array}\]

OpenStudy (unklerhaukus):

\[ \begin{pmatrix} 1&2&3 \\4&5&6 \\7&8&9 \end{pmatrix}, \begin{bmatrix} 10&11&12 \\13&14&15 \\16&17&18 \end{bmatrix}, \begin{vmatrix} 19&20&21 \\22&23&24 \\25&26&27 \end{vmatrix} \] ``` \[ \begin{pmatrix} 1&2&3 \\4&5&6 \\7&8&9 \end{pmatrix}, \begin{bmatrix} 10&11&12 \\13&14&15 \\16&17&18 \end{bmatrix}, \begin{vmatrix} 19&20&21 \\22&23&24 \\25&26&27 \end{vmatrix} \]